How many grams of Na2CO3 (FW = 106) are needed to prepare:

a) 250 mL of 6.0 M solution

b) 250 g of a 10% solution
Answer #1

a.

volume of solution = 250ml

molarity of solution = 6M

FW of Na2CO3 = 106g/mole

molarity = mass of Na2CO3*1000/FW of Na2CO3* volume of solution in ml

          6   = mass of Na2CO3*1000/106*250

mass of Na2CO3 = 6*106*250/1000    = 159g >>>>>answer

b.10% Na2CO3 means 10g of Na2CO3 in present in 100g of solution

100g of solution contains 10g of Na2CO3

250g of solution contains = 10*250/100   = 25g of Na2CO3 >>>>answer